# Asymptotic behaviors of subcritical branching killed spectrally negative Lévy process

### v1

@@ 3. Conditioned limit theorems for Lévy processes
×$\quad\times\mathbf{E}_{x}\left(\xi_{k};\tau_{0}^{-}>k,\xi_{k}\leq t^{(1-\epsilon)/2},\tau_{t^{1/2-\epsilon}}^{S,+}=k\right)$
@@
(3.34) $I_{4}\leq\frac{2R(x)\left(1+\frac{c_{14}(\epsilon)}{t^{\epsilon/2-\epsilon^{p}}}\right)\left(\mathcal{R}(y)+\frac{c_{14}(\epsilon)}{t^{\epsilon/2}}\right)}{\sigma\sqrt{2\pi t}}\leq\frac{2R(x)\mathcal{R}(y)}{\sigma\sqrt{2\pi t}}+\frac{c_{15}(\epsilon)(1+x)}{t^{\frac{1}{2}+\epsilon/4}}.$
(v) Lower bound of $I_{4}$. Repeating the proof of [6, (7.40)], we get that there exist positive constants $t_{9}(\epsilon)$, $c_{16}(\epsilon)$ and $c_{17}(\epsilon)$ such that for $t>t_{9}(\epsilon)$,
@@
(3.36) $\geq\frac{2R(x)\mathcal{R}(y)}{\sigma\sqrt{2\pi t}}-\frac{c_{17}(\epsilon)(1+x)}{t^{\frac{1}{2}+\epsilon/4}}.$
Set $\epsilon_{0}:=\min\{\delta/(2(5+2\delta)),\epsilon_{1}\}$, $\varepsilon:=\epsilon/4$, $\varepsilon_{0}:=\epsilon_{0}/4$ and $T_{0}(\epsilon):=\max\{t_{i}(\epsilon):i=1,2,8,9\}$. Using the fact that there exists $c_{18}>0$ such that $R(x)\leq c_{18}(1+x)$, and combining (3.12), (3.13), (3.14), (3.34) and (3.35), we arrive at the conclusion of the lemma. ∎
@@
For any $t>0$ and any bounded Borel functions $g,h:\mathbb{R}\mapsto\mathbb{R}_{+}$, we have
(3.37) $\int_{\mathbb{R}_{+}}h(x)\mathbf{E}_{x}\left[g(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right]\mathrm{d}x=\int_{\mathbb{R}_{+}}g(y)\mathbf{E}_{y}\left[h(\widehat{\xi}_{t})1_{\{\widehat{\tau}_{0}^{-}>t\}}\right]\mathrm{d}y$
@@
(3.38) $\int_{\mathbb{R}}h(x)\mathbf{E}_{x}\left[g(\xi_{t})1_{\{\tau_{0}^{-}\leq t\}}\right]\mathrm{d}x=\int_{\mathbb{R}}g(y)\mathbf{E}_{y}\left[h(\widehat{\xi}_{t})1_{\{\widehat{\tau}_{0}^{-}\leq t\}}\right]\mathrm{d}y.$
@@
(3.39) $\int_{\mathbb{R}_{+}}h(x)\mathbf{E}_{x}\left[g(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right]\mathrm{d}x=\int_{\mathbb{R}_{+}}h(x)\mathbf{E}\left[g(x+\xi_{t})1_{\{\tau_{-x}^{-}>t\}}\right]\mathrm{d}x$
(3.40) $=\int_{\mathbb{R}_{+}}h(x)\mathbf{E}\left[g(x+\xi_{t}),\inf_{s\leq t}\xi_{s}>-x\right]\mathrm{d}x=\int_{\mathbb{R}_{+}}h(x)\mathbf{E}\left[g(x+\xi_{t}),\inf_{s\leq t}\xi_{t-s}>-x\right]\mathrm{d}x$
(3.41) $=\int_{\mathbb{R}_{+}}g(y)\mathbf{E}\left[h(y-\xi_{t}),\inf_{s\leq t}(\xi_{t-s}-\xi_{t})>-y\right]\mathrm{d}y$
(3.42) $=\int_{\mathbb{R}_{+}}g(y)\mathbf{E}\left[h(y+\widehat{\xi}_{t}),\inf_{s\leq t}\widehat{\xi}_{s}>-y\right]\mathrm{d}y=\int_{\mathbb{R}_{+}}g(y)\mathbf{E}_{y}\left[h(\widehat{\xi}_{t})1_{\{\widehat{\tau}_{0}^{-}>t\}}\right]\mathrm{d}y,$
@@
(3.45) $\int_{\mathbb{R}}h(x)\mathbf{E}_{x}\left[g(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right]\mathrm{d}x=\int_{\mathbb{R}}g(y)\mathbf{E}_{y}\left[h(\widehat{\xi}_{t})1_{\{\widehat{\tau}_{0}^{-}>t\}}\right]\mathrm{d}y.$
@@
Before stating Theorem 3.5, we first introduce some necessary notation and definitions. Let $h_{1},h_{2}:\mathbb{R}\mapsto\mathbb{R}_{+}$ be Borel functions and $\varepsilon>0$. We say that $h_{1}$ $\varepsilon$-dominates $h_{2}$ and write $h_{2}\leq_{\varepsilon}h_{1}$ if
@@
For any $a>0$ and Borel function $h:\mathbb{R}\mapsto\mathbb{R}_{+}$, we define $I_{k,a}=[ka,(k+1)a]$ for $k\in\mathbb{Z}$ and
@@
(3.50) $\underline{h}_{a,-\varepsilon}\leq_{\varepsilon}\underline{h}_{a}\leq h\leq\bar{h}_{a}\leq_{\varepsilon}\bar{h}_{a,\varepsilon}\quad\mbox{on }\mathbb{R}.$
For more details about directly Riemann integrability, see [5, Section XI.1].
The following theorem will play an important role in this paper. We refer the reader to [7, Theorem 1.9] for an analogous result for random walks.
@@
Assume that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1), (H2) and $\mathbf{E}_{0}\left[\xi_{1}\right]<0$. Let $f:\mathbb{R}\mapsto\mathbb{R}_{+}$ be a Borel function, which is not 0 almost everywhere on $\mathbb{R}_{+}$, such that $f(x)e^{-\lambda_{*}x}(1+|x|)$ is directly Riemann integrable. Then for any $x>0$, it holds that
@@
Recall that, when $\mathbf{E}_{0}\left[\xi_{1}\right]<0$, $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{\lambda_{*}})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process with Laplace exponent $\Psi_{\lambda_{*}}(\lambda)=\Psi(\lambda+\lambda_{*})-\Psi(\lambda_{*})$ and that $\Psi_{\lambda_{*}}^{\prime}(0+)=\Psi^{\prime}(\lambda_{*})=0$. Using (1.24), we get that
@@
Assume that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1), (H2) and $\mathbf{E}_{0}\left[\xi_{1}\right]<0$. Let $h:\mathbb{R}\mapsto\mathbb{R}_{+}$ be a Borel function, which is not 0 almost everywhere on $\mathbb{R}_{+}$, such that $h(x)(1+|x|)$ is directly Riemann integrable. Then for any $x>0$, it holds that
@@
Taking $h(x)=f(x)e^{-\lambda_{*}x}$ in Theorem 3.6 and using (3.52), we immediately get the conclusion of Theorem 3.5.
In the next four lemmas, we provide some upper and lower bounds for $\mathbf{E}_{x}^{\lambda_{*}}\left(h(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right)$. In the remainder of this section, $\varepsilon_{0}$ will be the constant in Lemma 3.3. Recall that $\rho(\cdot)$ stands for the Rayleigh density.
@@
Assume that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1), (H2) and $\mathbf{E}_{0}\left[\xi_{1}\right]<0$. Then one can find a constant $C_{5}>0$ with the property that for any $\varepsilon\in(0,\varepsilon_{0})$ there exist positive constants $T_{1}(\varepsilon)$ and $C_{6}(\varepsilon)$ such that for any $x>0$, $t>T_{1}(\varepsilon)$ and any integrable functions $h,H:\mathbb{R}\mapsto\mathbb{R}_{+}$ satisfying $h\leq_{\varepsilon}H$,
@@
Fix $\varepsilon\in(0,\varepsilon_{0})$ and let $h,H:\mathbb{R}\mapsto\mathbb{R}_{+}$ be integrable functions satisfying $h\leq_{\varepsilon}H$. Fix $t\geq 1$ and set $m=[\varepsilon t]$. By the Markov property,
@@
Define a random walk $(S_{n})_{n\geq 0}$ by $S_{n}:=\xi_{n}$, $n\in\mathbb{N}$. Since $h1_{[0,\infty)}\leq_{\varepsilon}H1_{[-\varepsilon,\infty)}$, it follows from [7, Theorem 2.7] that there exist constants $c_{1}$ (independent of $\varepsilon)$ and $c_{2}(\varepsilon)$ such that for any $n\geq 1$,
@@
(3.62) $\mathbf{E}_{x}^{\lambda_{*}}\left(h(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right)$ $\leq\int_{\mathbb{R}_{+}}\left[\frac{1+c_{1}\varepsilon}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\int_{\mathbb{R}}H(z)1_{\{z\geq-\varepsilon\}}\phi\left(\frac{z-y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\right)\mathrm{d}z\right]$
×$\quad\times\mathbf{P}_{x}^{\lambda_{*}}\left(\xi_{t-m}\in\mathrm{d}y,\tau_{0}^{-}>t-m\right)$
@@
$A_{1}(x)$ $=\int_{\mathbb{R}_{+}}\left[\frac{1+c_{1}\varepsilon}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\int_{\mathbb{R}}H(z)1_{\{z\geq-\varepsilon\}}\phi\left(\frac{z-\sqrt{\Psi^{\prime\prime}(\lambda_{*})(t-m)}u}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\right)\mathrm{d}z\right]$
×$\quad\times\mathbf{P}_{x}^{\lambda_{*}}\left(\frac{\xi_{t-m}}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})(t-m)}}\in\mathrm{d}u,\tau_{0}^{-}>t-m\right)$
@@
According to [6, Lemma 3.3], for any $v\in(0,\frac{1}{2}]$ and $s\geq 0$, it holds that
@@
(3.76) $\leq\frac{(1+c_{1}\varepsilon)}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})t}}\int^{\infty}_{-\varepsilon}H(w)\left[\sqrt{\frac{t-m}{t}}\rho\left(\frac{w}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})t}}\right)+\sqrt{\frac{m}{t}}e^{-\frac{w^{2}}{2\Psi^{\prime\prime}(\lambda_{*})t}}\right]\mathrm{d}w.$
@@
Assume that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1), (H2) and $\mathbf{E}_{0}\left[\xi_{1}\right]<0$. Then one can find a constant $C_{7}>0$ with the property that for any $\varepsilon\in(0,\varepsilon_{0})$ there exist positive constants $T_{2}(\varepsilon)$ and $C_{8}(\varepsilon)$ such that for any $x>0$, $t>T_{2}(\varepsilon)$ and any Borel functions $h,H:{\mathbb{R}}\mapsto\mathbb{R}_{+}$ satisfying $h\leq_{\varepsilon}H$ and $\int_{\mathbb{R}_{+}}H(z-\varepsilon)(1+z)\mathrm{d}z<\infty$,
@@
Fix $\varepsilon\in(0,\varepsilon_{0})$ and let $h,H:\mathbb{R}\mapsto\mathbb{R}_{+}$ satisfying $h\leq_{\varepsilon}H$. For any $z\in\mathbb{R}$, we define
@@
Assume that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1), (H2) and $\mathbf{E}_{0}(\xi_{1})<0$. Then one can find positive constants $C_{9}$ and $q$ with the property that for any $\varepsilon\in(0,\varepsilon_{0})$ there exist positive constants $T_{3}(\varepsilon)$ and $C_{10}(\varepsilon)$ such that for any $x>0$, $t>T_{3}(\varepsilon)$ and any integrable functions $h,H,g:\mathbb{R}\mapsto\mathbb{R}_{+}$ satisfying $g\leq_{\varepsilon}h\leq_{\varepsilon}H$,
$\mathbf{E}_{x}^{\lambda_{*}}\left(h(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right)\geq$ $\frac{2R^{*}(x)}{\sqrt{2\pi}\Psi^{\prime\prime}(\lambda_{*})t}\int_{\mathbb{R}_{+}}\left[g(w)1_{\{w\geq\varepsilon\}}-C_{9}\varepsilon h(w)\right]\rho\left(\frac{w}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})t}}\right)\mathrm{d}w$
@@
Fix $\varepsilon\in(0,\varepsilon_{0})$ and let $h,H,g:\mathbb{R}\mapsto\mathbb{R}_{+}$ be integrable functions satisfying $g\leq_{\varepsilon}h\leq_{\varepsilon}H$. Then $g1_{[\varepsilon,\infty)}\leq_{\varepsilon}h1_{[0,\infty)}\leq_{\varepsilon}H1_{[-\varepsilon,\infty)}$. Fix $t\geq 1$ and set $m=[\varepsilon t]$. By the Markov property,
@@
Step 1. In this step, we give a lower bound for $I_{1}(t)$. By [7, Theorem 2.7], there exist positive constants $c_{1}$ (independent of $\varepsilon$) and $c_{2}(\varepsilon)$ such that for any $m\geq 1$,
(3.108) $\mathbf{E}_{y}^{\lambda_{*}}\left(h(\xi_{m})1_{\{\xi_{m}\geq 0\}}\right)$ $\geq\frac{1}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\int_{\mathbb{R}}\left[g(z)1_{\{z\geq\varepsilon\}}-c_{1}\varepsilon h(z)1_{\{z\geq 0\}}\right]\phi\left(\frac{z-y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\right)\mathrm{d}z$
@@
$\frac{1}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\int_{\mathbb{R}_{+}}\left[\int_{\mathbb{R}}g(z)1_{\{z\geq\varepsilon\}}\phi\left(\frac{z-y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\right)\mathrm{d}z\right]\mathbf{P}_{x}^{\lambda_{*}}\left(\xi_{t-m}\in\mathrm{d}y,\tau_{0}^{-}>t-m\right)$
@@
$\frac{c_{1}\varepsilon}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\int_{\mathbb{R}_{+}}\left[\int_{\mathbb{R}}h(z)1_{\{z\geq 0\}}\phi\left(\frac{z-y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\right)\mathrm{d}z\right]\mathbf{P}_{x}^{\lambda_{*}}\left(\xi_{t-m}\in\mathrm{d}y,\tau_{0}^{-}>t-m\right)$

### v2

@@ 3. Conditioned limit theorems for Lévy processes
$\quad\times\mathbf{E}_{x}\left(\xi_{k};\tau_{0}^{-}>k,\xi_{k}\leq t^{(1-\epsilon)/2},\tau_{t^{1/2-\epsilon}}^{S,+}=k\right)$
@@
(3.34) $I_{4}\leq\frac{2R(x)\left(1+\frac{c_{14}(\epsilon)}{t^{\epsilon/2-\epsilon^{p}}}\right)\left(\mathcal{R}(y)+\frac{c_{14}(\epsilon)}{t^{\epsilon/2}}\right)}{\sigma\sqrt{2\pi t}}\leq\frac{2R(x)\mathcal{R}(y)}{\sigma\sqrt{2\pi t}}+\frac{c_{15}(\epsilon)(1+x)}{t^{\frac{1}{2}+\epsilon/8}},$
where in the last inequality we use the fact that $R(x)\leq c(1+x)$ for some constant $c>0$.
(v) Lower bound of $I_{4}$. Repeating the proof of [7, (7.40)], we get that there exist positive constants $t_{9}(\epsilon)$, $c_{16}(\epsilon)$ and $c_{17}(\epsilon)$ such that for $t>t_{9}(\epsilon)$,
@@
(3.36) $\geq\frac{2R(x)\mathcal{R}(y)}{\sigma\sqrt{2\pi t}}-\frac{c_{17}(\epsilon)(1+x)}{t^{\frac{1}{2}+\epsilon/8}}.$
Set $\epsilon_{0}:=\min\{\delta/(4(5+2\delta)),\epsilon_{1}\}$, $\varepsilon:=\epsilon/8$, $\varepsilon_{0}:=\epsilon_{0}/8$ and $T_{0}(\epsilon):=\max\{t_{i}(\epsilon):i=1,2,8,9\}$. Using the fact that there exists $c_{18}>0$ such that $R(x)\leq c_{18}(1+x)$, and combining (3.12), (3.13), (3.14), (3.34) and (3.35), we arrive at the conclusion of the lemma. ∎
@@
For any $t>0$ and any bounded Borel functions $g,h:\mathbb{R}\to\mathbb{R}_{+}$, we have
(3.37) $\int_{\mathbb{R}_{+}}h(x)\mathbf{E}_{x}\left(g(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right)\mathrm{d}x=\int_{\mathbb{R}_{+}}g(y)\mathbf{E}_{y}\left(h(\widehat{\xi}_{t})1_{\{\widehat{\tau}_{0}^{-}>t\}}\right)\mathrm{d}y$
@@
(3.38) $\int_{\mathbb{R}}h(x)\mathbf{E}_{x}\left(g(\xi_{t})1_{\{\tau_{0}^{-}\leq t\}}\right)\mathrm{d}x=\int_{\mathbb{R}}g(y)\mathbf{E}_{y}\left(h(\widehat{\xi}_{t})1_{\{\widehat{\tau}_{0}^{-}\leq t\}}\right)\mathrm{d}y.$
@@
(3.39) $\int_{\mathbb{R}_{+}}h(x)\mathbf{E}_{x}\left(g(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right)\mathrm{d}x=\int_{\mathbb{R}_{+}}h(x)\mathbf{E}\left(g(x+\xi_{t})1_{\{\tau_{-x}^{-}>t\}}\right)\mathrm{d}x$
(3.40) $=\int_{\mathbb{R}_{+}}h(x)\mathbf{E}\left(g(x+\xi_{t}),\inf_{s\leq t}\xi_{s}>-x\right)\mathrm{d}x=\int_{\mathbb{R}_{+}}h(x)\mathbf{E}\left(g(x+\xi_{t}),\inf_{s\leq t}\xi_{t-s}>-x\right)\mathrm{d}x$
(3.41) $=\int_{\mathbb{R}_{+}}g(y)\mathbf{E}\left(h(y-\xi_{t}),\inf_{s\leq t}(\xi_{t-s}-\xi_{t})>-y\right)\mathrm{d}y$
(3.42) $=\int_{\mathbb{R}_{+}}g(y)\mathbf{E}\left(h(y+\widehat{\xi}_{t}),\inf_{s\leq t}\widehat{\xi}_{s}>-y\right)\mathrm{d}y=\int_{\mathbb{R}_{+}}g(y)\mathbf{E}_{y}\left(h(\widehat{\xi}_{t})1_{\{\widehat{\tau}_{0}^{-}>t\}}\right)\mathrm{d}y,$
@@
(3.45) $\int_{\mathbb{R}}h(x)\mathbf{E}_{x}\left(g(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right)\mathrm{d}x=\int_{\mathbb{R}}g(y)\mathbf{E}_{y}\left(h(\widehat{\xi}_{t})1_{\{\widehat{\tau}_{0}^{-}>t\}}\right)\mathrm{d}y.$
@@
Before stating Theorem 3.5, we first introduce some necessary notation and definitions. Let $h_{1},h_{2}:\mathbb{R}\to\mathbb{R}_{+}$ be Borel functions and $\varepsilon>0$. We say that $h_{1}$ $\varepsilon$-dominates $h_{2}$ and write $h_{2}\leq_{\varepsilon}h_{1}$ if
@@
For any $a>0$ and Borel function $h:\mathbb{R}\to\mathbb{R}_{+}$, we define $I_{k,a}=[ka,(k+1)a]$ for $k\in\mathbb{Z}$ and
@@
[…]
For more details about directly Riemann integrability, see [6, Section XI.1].
The following theorem will play an important role in this paper. We refer the reader to [8, Theorem 1.9] for an analogous result for random walks.
@@
Assume that $\xi$ is a Lévy process satisfying (H1), (H2), (H3) and $\mathbf{E}_{0}\left(\xi_{1}\right)<0$. Let $f:\mathbb{R}\to\mathbb{R}_{+}$ be a Borel function, which is not 0 almost everywhere on $\mathbb{R}_{+}$, such that $f(x)e^{-\lambda_{*}x}(1+|x|)$ is directly Riemann integrable. Then for any $x>0$, it holds that
@@
Recall that when $\mathbf{E}_{0}\left(\xi_{1}\right)<0$ and (H3) holds, $\xi^{(\lambda_{*})}$ is a Lévy process with Laplace exponent $\Psi_{\lambda_{*}}(\lambda)=\Psi(\lambda+\lambda_{*})-\Psi(\lambda_{*})$ and that $\Psi_{\lambda_{*}}^{\prime}(0+)=\Psi^{\prime}(\lambda_{*})=0$. Using (1.21), we get that
@@
Assume that $\xi$ is a Lévy process satisfying (H1), (H2), (H3) and $\mathbf{E}_{0}\left(\xi_{1}\right)<0$. Let $h:\mathbb{R}\to\mathbb{R}_{+}$ be a Borel function, which is not 0 almost everywhere on $\mathbb{R}_{+}$, such that $h(x)(1+|x|)$ is directly Riemann integrable. Then for any $x>0$, it holds that
@@
The difference between the asymptotic behavior in Theorem 3.6 (with $t^{-3/2}$ decay) and that in Lemma 3.1 (with $t^{-1/2}$ decay) arises from the fact that Theorem 3.6 is a conditioned limit theorem for the process $\left(\xi_{t}\right)_{t\geq 0}$ itself, whereas Lemma 3.1 is a conditioned limit result for the normalized process $\left(\frac{\xi_{t}}{\sigma\sqrt{t}}\right)_{t\geq 0}$.
Proof of Theorem 3.5: Taking $h(x)=f(x)e^{-\lambda_{*}x}$ in Theorem 3.6 and using (3.52), we immediately get the conclusion of Theorem 3.5. ∎
In the next four lemmas, we provide some upper and lower bounds for $\mathbf{E}_{x}^{\lambda_{*}}\left(h(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right)$. In the remainder of this section, $\varepsilon_{0}$ will be the constant in Lemma $3.$ Recall that $\rho(\cdot)$ stands for the Rayleigh density.
@@
Assume that $\xi$ is a Lévy process satisfying (H1), (H2), (H3) and $\mathbf{E}_{0}\left(\xi_{1}\right)<0$. Then one can find a constant $C_{5}>0$ with the property that for any $\varepsilon\in(0,\varepsilon_{0})$ there exist positive constants $T_{1}(\varepsilon)$ and $C_{6}(\varepsilon)$ such that for any $x>0$, $t>T_{1}(\varepsilon)$ and any integrable functions $h,H:\mathbb{R}\to\mathbb{R}_{+}$ satisfying $h\leq_{\varepsilon}H$,
@@
Fix $\varepsilon\in(0,\varepsilon_{0})$ and let $h,H:\mathbb{R}\to\mathbb{R}_{+}$ be integrable functions satisfying $h\leq_{\varepsilon}H$. Fix $t\geq 1$ and set $m=[\varepsilon t]$. By the Markov property,
@@
Define a random walk $(S_{n})_{n\geq 0}$ by $S_{n}:=\xi_{n}$, $n\in\mathbb{N}$. Since $h1_{[0,\infty)}\leq_{\varepsilon}H1_{[-\varepsilon,\infty)}$, it follows from [8, Theorem 2.7] that there exist constants $c_{1}$ (independent of $\varepsilon)$ and $c_{2}(\varepsilon)$ such that for any $n\geq 1$,
@@
(3.62) $\mathbf{E}_{x}^{\lambda_{*}}\left(h(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right)$ $\leq\int_{\mathbb{R}_{+}}\left(\frac{1+c_{1}\varepsilon}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\int_{\mathbb{R}}H(z)1_{\{z\geq-\varepsilon\}}\phi\left(\frac{z-y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\right)\mathrm{d}z\right)$
$\quad\times\mathbf{P}_{x}^{\lambda_{*}}\left(\xi_{t-m}\in\mathrm{d}y,\tau_{0}^{-}>t-m\right)$
@@
$A_{1}(x)$ $=\int_{\mathbb{R}_{+}}\left(\frac{1+c_{1}\varepsilon}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\int_{\mathbb{R}}H(z)1_{\{z\geq-\varepsilon\}}\phi\left(\frac{z-\sqrt{\Psi^{\prime\prime}(\lambda_{*})(t-m)}u}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\right)\mathrm{d}z\right)$
$\quad\times\mathbf{P}_{x}^{\lambda_{*}}\left(\frac{\xi_{t-m}}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})(t-m)}}\in\mathrm{d}u,\tau_{0}^{-}>t-m\right)$
@@
According to [7, Lemma 3.3], for any $v\in(0,\frac{1}{2}]$ and $s\geq 0$, it holds that
@@
(3.76) $\leq\frac{(1+c_{1}\varepsilon)}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})t}}\int^{\infty}_{-\varepsilon}H(w)\left(\sqrt{\frac{t-m}{t}}\rho\left(\frac{w}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})t}}\right)+\sqrt{\frac{m}{t}}e^{-\frac{w^{2}}{2\Psi^{\prime\prime}(\lambda_{*})t}}\right)\mathrm{d}w.$
@@
Assume that $\xi$ is a Lévy process satisfying (H1), (H2), (H3) and $\mathbf{E}_{0}\left(\xi_{1}\right)<0$. Then one can find a constant $C_{7}>0$ with the property that for any $\varepsilon\in(0,\varepsilon_{0})$ there exist positive constants $T_{2}(\varepsilon)$ and $C_{8}(\varepsilon)$ such that for any $x>0$, $t>T_{2}(\varepsilon)$ and any Borel functions $h,H:{\mathbb{R}}\to\mathbb{R}_{+}$ satisfying $h\leq_{\varepsilon}H$ and $\int_{\mathbb{R}_{+}}H(z-\varepsilon)(1+z)\mathrm{d}z<\infty$,
@@
Fix $\varepsilon\in(0,\varepsilon_{0})$ and let $h,H:\mathbb{R}\to\mathbb{R}_{+}$ satisfying $h\leq_{\varepsilon}H$. For any $z\in\mathbb{R}$, we define
@@
Assume that $\xi$ is a Lévy process satisfying (H1), (H2), (H3) and $\mathbf{E}_{0}(\xi_{1})<0$. Then one can find positive constants $C_{9}$ and $q$ with the property that for any $\varepsilon\in(0,\varepsilon_{0})$ there exist positive constants $T_{3}(\varepsilon)$ and $C_{10}(\varepsilon)$ such that for any $x>0$, $t>T_{3}(\varepsilon)$ and any integrable functions $h,H,g:\mathbb{R}\to\mathbb{R}_{+}$ satisfying $g\leq_{\varepsilon}h\leq_{\varepsilon}H$,
$\mathbf{E}_{x}^{\lambda_{*}}\left(h(\xi_{t})1_{\{\tau_{0}^{-}>t\}}\right)\geq$ $\frac{2R^{*}(x)}{\sqrt{2\pi}\Psi^{\prime\prime}(\lambda_{*})t}\int_{\mathbb{R}_{+}}\left(g(w)1_{\{w\geq\varepsilon\}}-C_{9}\varepsilon h(w)\right)\rho\left(\frac{w}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})t}}\right)\mathrm{d}w$
@@
Fix $\varepsilon\in(0,\varepsilon_{0})$ and let $h,H,g:\mathbb{R}\to\mathbb{R}_{+}$ be integrable functions satisfying $g\leq_{\varepsilon}h\leq_{\varepsilon}H$. Then $g1_{[\varepsilon,\infty)}\leq_{\varepsilon}h1_{[0,\infty)}\leq_{\varepsilon}H1_{[-\varepsilon,\infty)}$. Fix $t\geq 1$ and set $m=[\varepsilon t]$. By the Markov property,
@@
Step 1. In this step, we give a lower bound for $I_{1}(t)$. By [8, Theorem 2.7], there exist positive constants $c_{1}$ (independent of $\varepsilon$) and $c_{2}(\varepsilon)$ such that for any $m\geq 1$,
(3.108) $\mathbf{E}_{y}^{\lambda_{*}}\left(h(\xi_{m})1_{\{\xi_{m}\geq 0\}}\right)$ $\geq\frac{1}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\int_{\mathbb{R}}\left(g(z)1_{\{z\geq\varepsilon\}}-c_{1}\varepsilon h(z)1_{\{z\geq 0\}}\right)\phi\left(\frac{z-y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\right)\mathrm{d}z$
@@
$\frac{1}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\int_{\mathbb{R}_{+}}\left(\int_{\mathbb{R}}g(z)1_{\{z\geq\varepsilon\}}\phi\left(\frac{z-y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\right)\mathrm{d}z\right)\mathbf{P}_{x}^{\lambda_{*}}\left(\xi_{t-m}\in\mathrm{d}y,\tau_{0}^{-}>t-m\right)$
@@
$\frac{c_{1}\varepsilon}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\int_{\mathbb{R}_{+}}\left(\int_{\mathbb{R}}h(z)1_{\{z\geq 0\}}\phi\left(\frac{z-y}{\sqrt{\Psi^{\prime\prime}(\lambda_{*})m}}\right)\mathrm{d}z\right)\mathbf{P}_{x}^{\lambda_{*}}\left(\xi_{t-m}\in\mathrm{d}y,\tau_{0}^{-}>t-m\right)$